cle, a researcher assertsthat here have been changes in spontaneous ou door acti ties such as bike ding, s 1 ming, and ouch oIncludes / does not include seems / does not seem
-za za +/-za/2
reject / do not reject insufficient / sufficient
Answer #1

Basic data:

\hat{p}_1 = Sample proportion of 7-11 year olds who rode their bikes at least 6 time in 1994 = 0.72 and 1 - \hat{p}_1 = 0.28

n1 = 2000, x1 = 2000 * 0.72 = 1440

\hat{p}_2 = Sample proportion of 7-11 year olds who rode their bikes at least 6 time in 2003 = 0.45, and 1 - \hat{p}_2 = 0.55

n2 = 2500, x2 = 2500 * 0.45 = 1125

\hat{p} = Overall Proportion = (1440 + 1125) / (2000 + 2500) = 2565//4500 = 0.57

1 - \hat{p} = 0.43

\alpha = 0.02

Zcritical value (2 tailed) = -2.33and +2.333

_________________________________________________________________

(a) 98% Confidence Interval

The Confidence Interval is given by (P1- \hat{p_2} ) \pm ME, where

(P1- \hat{p_2} ) = 0.72 – 0.45 = 0.27

PMI-P1) . P2(1-p2) 0.72 0.28 0.45 0.55 NIE Zcritical* 2.33* 1 12 2500 0,0329

The Lower Limit = 0.27 - 0.0329 = 0.2371

The Upper Limit = 0.27 + 0.0329 = 0.3029

The 98% Confidence Interval is 0.2371 < p1 - p2 < 0.3029

Since the interval (0.2371,0.3029) does not contain 0, it is not plausible that the proportion of 7 to 11 year olds who rode their bikes 6 times in 2003 is the same as in 1994.

__________________________________________________________________________

(b) 2 tailed test for difference in proportions

The Hypothesis: Option C

H0: p1 - p2 = 0

Ha: p1 - p2 \neq 0

The critical values are Z = -2.33,2.33

The Test Statistic:

pl - p2 0.72-0.45 = 19.149 2000 2500

Z = 19.149

Reject the Null Hypothesis. There is sufficient evidence that the proportion of 7 to 11 year olds who rode their bikes at least 6 times in 2003 is different from that in 1994.

Are the results contradictory: Option B: No, because the results from both parts indicate that the proportions for the two years are different.